# REALM-Bench: A Real-World Planning Benchmark for LLMs and Multi-Agent Systems

### v1

@@ 4.1. P1: Campus Single-Tour Navigation
    Environment: A known map with a finite set of locations.
    Goal: Visit all designated waypoints within a given timeframe.
    Constraints: Opening hours of each location, each location at least 30 minutes, and must be completed before 5 PM.
    Optimization Metric: Shortest path (time or distance).
A meta-plan, shown in Table 1, provides the high-level structure and constraints for the problem. This meta-plan serves as input to specialized solvers, such as dynamic programming or Monte Carlo algorithms, which then generate detailed, executable workflows. The process transforms abstract planning requirements into concrete, implementable sequences of actions while respecting all specified constraints and optimization objectives.
4.2. P2: Multi-Group Campus Tour
Problem Statement: Multiple groups of visitors require guided tours in different locations on a university campus, with optimized scheduling of multiple tour guides. This problem shares the same metrics as P1.
Problem Specification:
    Multiple agents (tour guides) must coordinate to serve different groups of visitors.
    Each group has predefined preferences and time constraints.
    Agents must follow non-overlapping paths while minimizing idle time.
    Location visiting hours must be observed.
Details are provided in Table 2.
4.3. P3: Urban Ride-Sharing (URS)
Problem Statement: Optimize real-time ride assignments for multiple vehicles and passengers in an urban environment, balancing efficiency, fuel use, and service quality.
Problem Specification:
    City Map: A graph $G=(V,E)$ where locations $V$ and roads $E$ have distances and travel times.
    Ride Requests: A set of requests $R$, each defined by:
    • Passenger ID, pickup/drop-off locations, time windows.
    Vehicles: A set of available vehicles $K$, each with:
    • Location, battery/fuel level, passenger capacity and speed.
See Table 3 for detailed specifications.
P4: Urban Ride Sharing (URS) with Disruptions
Problem Specification: Details are provided in Table 4, including two disruption scenarios: airport route traffic delay and local road closure. Other possible disruptions could involve cancelation of passengers or late arrivals.
@@ 4.3. P3: Urban Ride-Sharing (URS)
Vehicles: Three vehicles $V=\{v_{1},v_{2},v_{3}\}$ Capacity: 2 passengers each - Initial vehicle location: City center - Operating hours: all day Passengers: Five passengers $P=\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$ $p_{1}$: Airport by 8:30 AM - $p_{2}$: Airport by 9:00 AM - $p_{3}$: Airport by 9:30 AM - $p_{4}$: Airport by 9:45 AM - $p_{5}$: Airport by 9:45 AM Travel Times: - City center to pickup locations: 15-30 minutes - Pickup locations to airport: 45-60 minutes - Between pickup locations: 20-30 minutes Disruptions: - Airport route traffic delay - Certain local road closure Objectives: - Minimize total vehicle travel time - Meet all passenger deadlines
P5: Wedding Reunion
Problem Specification: Table 5 presents a coordinated wedding event travel problem. Several friends arrive at different times and locations before a 3:00 PM wedding photo session. The challenge includes managing two vehicles for airport pick-ups (aimed at those who cannot drive or wish to cut costs) and completing critical errands, such as collecting the wedding gift and retrieving formal attire from the tailor. All activities must be scheduled to ensure that everyone arrives at the wedding venue before the photo time.
This problem introduces more constraints than the URS problems in P3 and P4, and it also lays the groundwork for a more challenging disruption case discussed in P8.
@@
Metrics: - On-time performance: Must arrive at the venue for 3:00 PM photos. Locations: Four locations: $V=\{B,G,T,W\}$, where $B$ is Boston Airport, $G$ is Gift shop, $T$ is Tailor shop, and $W$ is Wedding venue. Travel time: (minutes) ${}B\text{-}G:45,~{}B\text{-}T:30,~{}B\text{-}W:40,~{}G\text{-}T:20,~{}G\text{-}W:25,~{}T\text{-}W:15$. Arrival Times: - Alex: At $B$ at 11:00 AM from Chicago (need a ride) - Jamie: At $B$ at 12:30 PM from Atlanta (need a ride) - Pat: At $W$ at 12:00 PM driving from NYC (has 5-seater car) Required Tasks: - Gift collection from $G$ (after 12:00 PM) - Clothes pickup from $T$ (by 2:00 PM) - Photos at $W$ (3:00 PM sharp) Available Resources: - One car (5-seater) with Pat, available after he is Boston - Local friend Chris (5-seater) available after 1:30 PM at $W$ Scheduling Constraints: - All tasks must complete before 3:00 PM photo time - Gift store opens at 12:00 PM - Tailor closes at 2:00 PM - Two cars must accommodate all transport needs
P6: Thanksgiving Dinner
Consider a Thanksgiving dinner scenario in which a family of five must return home in a Boston suburb for dinner. The problem involves coordinating departure times, managing travel logistics (including possible traffic delays), and ensuring timely arrival. Table 6 formalizes these challenges as a sequential planning problem.
This scenario also lays the groundwork for a more advanced disruption case, which has proven difficult for standalone LLMs, as discussed in P9.
Problem Specification:
Setup:
  • @itemi
    Mom (Sarah) hosts dinner at 6:00 PM in Boston.
  • @itemi
    Family arrivals:
    • Dad (James) from San Francisco, lands at 1:00 PM ET.
    • Sister (Emily) from Chicago lands at 2:30 PM.
    • Brother (Michael) driving from NY arrives at 3:00 PM.
  • @itemi
    Grandma, who is healthy, needs to pick up nearby.
Constraints:
  • @itemi
    James must rent a car post-landing.
  • @itemi
    Emily needs an airport pickup (no alternatives).
  • @itemi
    Turkey requires 4 hours to cook; someone must be home once it’s in the oven for safety.
  • @itemi
    Side dishes need 2 hours of preparation.
  • @itemi
    Travel times:
    • Home to BOS Airport: 1 hour.
    • BOS Airport to Grandma’s: 1 hour.
    • Home to Grandma’s: 30 minutes.
Key Planning Questions:
  1. 1.
    When should cooking start?
  2. 2.
    Who picks up Emily, and when?
  3. 3.
    When and by whom should Grandma be picked up?
@@
P7: Disaster Relief Logistics Problem
Problem Specification: Table 7 summarizes the problem.
@@
Network: $G=(V,E)$ where $V$ represents locations and $E$ represents routes Locations ($V$): - Supply nodes: CW (central warehouse), AP (airport, capacity: 5 tons) - Demand nodes: DC1, DC2, DC3 (distribution centers) - Critical nodes: H1, H2 (hospitals), FS (fuel station) Resources: $cap_{truck}$ $=4\text{ tons},\tau_{truck}=120\text{ min}$ $cap_{heli}$ $=1\text{ ton},\tau_{heli}=30\text{ min}$ Demand Requirements: - DCs: food/water delivery by $T_{dc}=20:00$ Hospitals: medicine within $\Delta T_{h}=6$ hours - Fuel Station: refuel by $T_{fs}=12:00$ Critical Deadlines: - Food/Water: All DCs by 8:00 PM - Medicine: Hospitals within 6 hours - Fuel: FS by 12:00 PM - Airport: Clear excess beyond 5 tons immediately Dynamic Disruptions: - Unpredictable donation arrivals - Road blockages requiring rerouting - Emergency hospital demands - Fuel shortage delays Key Planning Requirements: - Resource distribution scheduling - Transportation mode optimization - Delivery prioritization - Airport overflow management - Real-time disruption handling
P8: Wedding Reunion with Disruptions
@@
P9: Thanksgiving Dinner with Disruptions
@@
For example, consider James’s flight, which was originally scheduled to arrive in Boston at 1 PM but has been delayed to 4 PM. He learns of this new arrival time at 10 AM EST (the flight’s intended departure from SFO), offering a three-hour lead for adjustments. However, existing LLM-based solutions do not adapt the plan until 1 PM, thus squandering the opportunity to re-optimize the schedule.
P10: Global Supply Chain
Table 8. Summary of the Data Center Construction Problem Statement
Component Details
GPU Procurement & Shipments
Total GPU Target 100,000 units
Vendors NVIDIA: \$15k/unit, 200k units/quarter capacity, 20% maintenance risk (of unit price over 1 year) AMD: \$10k/unit, 150k units/quarter capacity, 50% maintenance risk (of unit price over 1 year)
Order Timing All orders placed on Day 1
Shipment Schedule Quarterly shipments; e.g., a 600,000 unit order results in 3 shipments of 200,000 units at the end of Q1, Q2, and Q3
Payment Terms 50% deposit at order (day 1), 50% at delivery (each quarter)
Natural Disaster Impacts
Risk Probabilities Earthquake: 10% per quarter; Typhoon: 10% per quarter (assessed over entire project)
Delay Impact Cascading 1-month delay on all subsequent shipments if a disaster occurs in a quarter
Cost Impact 30% price increase applied only to the affected quarter’s shipment
Shipment Expedition Impact Expedite one shipment by 1 month and all subsequent shipments are moved up by 1 month (lock stepped)
Cluster Construction & Infrastructure
Cluster Definition Each cluster requires 50,000 GPUs; clusters can be built concurrently once required GPUs are delivered
Networking Infrastructure (per cluster)
Duration 1 months
Cost \$25 million
Dependencies Can begin only after GPUs have arrived, power is ready, and cooling is operational
Power & Cooling (per cluster)
Requirements Power: 150 MW; Cooling: 1M gallons/day
Lead Time Options Regular: 2 months at \$30M per cluster Expedited: 1 months at \$75M per cluster (150% cost premium)
Testing & Certification (per cluster)
Normal Process 2 months, \$15 million
Expedited Process 1 month, \$45 million (150% cost premium)
Expedition Options Expedite power/cooling, networking (50% time reduction, 150% cost premium), and testing independently per cluster
Business Impact of Delays
Delay Costs \$50 million lost revenue + \$10 million additional operating costs per month of delay
Overall Objective
Objective Minimize the end-to-end construction cost (procurement, construction, maintenance, delay costs, expedition premiums) while meeting the 15-month deadline. If the 15-month not met, adding delay costs
Table 8 presents a comprehensive problem in data center GPU deployment that captures the complexity of large-scale infrastructure projects. The objective is to complete a 1 million GPU data center in 15 months while minimizing total costs. The problem encompasses procurement decisions between NVIDIA ($15$k/unit) and AMD ($10$k/unit) GPUs, where each vendor has different maintenance risks (20% vs 50% of unit price over one year) and quarterly shipment capacities.

### v2

@@ 3.3. P3: Urban Ride-Sharing Planner (URS-static, Table 3)
    Environment: Map $G=(V,E)$ with locations and roads
    Goal: Match vehicles to passenger requests efficiently
    Constraints: Vehicle capacity, battery/fuel levels, pickup/drop-off times
    Optimization Metric: Balance efficiency, fuel use, and service quality
3.4. P4: Urban Ride-Sharing with Disruptions (URS-dynamic, Table 4)
Problem Statement: Addition to P3 with dynamic disruptions requiring real-time reactive planning. Disruptions can include traffic delays, road closures, and passenger cancellations, et al.
@@ 3.4. P4: Urban Ride-Sharing with Disruptions (URS-dynamic, Table 4)
Vehicles: Three vehicles $V=\{v_{1},v_{2},v_{3}\}$ - Capacity: 2 passengers each - Initial vehicle location: City center - Operating hours: all day Passengers: Five passengers $P=\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$ - $p_{1}$: Airport by 8:30 AM; $p_{2}$: Airport by 9:00 AM - $p_{3}$: Airport by 9:30 AM; $p_{4}$: Airport by 9:45 AM - $p_{5}$: Airport by 9:45 AM Travel Times: - City center to pickup locations: 15-30 minutes - Pickup locations to airport: 45-60 minutes - Between pickup locations: 20-30 minutes Disruptions: - Airport route traffic delay - Certain local road closure Objectives: - Minimize total vehicle travel time - Meet all passenger deadlines
3.5. P5: Wedding Reunion Planner and Scheduler (WR-static, Table 5)
Problem Specification: Table 5 presents a coordinated travel problem for wedding events. Several friends arrive at different times and locations before a 3:00 PM wedding photo session. The challenge includes managing two vehicles for airport pick-ups (aimed at those who cannot drive or wish to cut costs) and completing critical errands, such as collecting the wedding gift and retrieving formal attire from the tailor. All activities must be scheduled to ensure that everyone arrives at the wedding venue before the photo session deadline.
This problem introduces more constraints than the URS problems in P3 and P4, and it also lays the groundwork for a more challenging disruption case depicted in P8.
@@ 3.5. P5: Wedding Reunion Planner and Scheduler (WR-static, Table 5)
Metrics: - On-time performance: Must be at the venue for 3:00 PM photos. Locations: Four locations: $V=\{B,G,T,W\}$, where $B$ is Boston Airport, $G$ is Gift shop, $T$ is Tailor shop, and $W$ is Wedding venue. Travel time: (minutes) $B\text{-}G:45,~B\text{-}T:30,~B\text{-}W:40,~G\text{-}T:20,~G\text{-}W:25,~T\text{-}W:15$. Arrival Times: - Alex: At $B$ at 11:00 AM from Chicago (need a ride) - Jamie: At $B$ at 12:30 PM from Atlanta (need a ride) - Pat: At $W$ at 12:00 PM driving from NYC (has 5-seater car) Required Tasks: - Gift collection from $G$ (after 12:00 PM) - Clothes pickup from $T$ (by 2:00 PM) - Photos at $W$ (3:00 PM sharp) Available Resources: - One car (5-seater) with Pat, available after he is Boston - Local friend Chris (5-seater) available after 1:30 PM at $W$ Scheduling Constraints: - All tasks must complete before 3:00 PM photo time - Gift store opens at 12:00 PM - Tailor closes at 2:00 PM - Two cars must accommodate all transport needs
3.6. P6: Thanksgiving Dinner Planner and Scheduler (TD-static, Table 6 )
Consider a Thanksgiving dinner scenario in which a family of five must return to their home in a Boston suburb for a 6 p.m. dinner. The problem involves coordinating departure times, managing travel logistics (including possible traffic delays), and ensuring timely arrival. Table 6 formalizes these challenges as a sequential planning problem.
This scenario also lays the foundation for a more advanced disruption case, which has proven difficult for standalone LLMs, as discussed in P9.
@@ 3.6. P6: Thanksgiving Dinner Planner and Scheduler (TD-static, Table 6 )
3.7. P7: Disaster Relief Logistics Planner and Scheduler (DL-static, Table 7 )
@@
Network: $G=(V,E)$ where $V$ represents locations and $E$ represents routes Locations ($V$): - Supply nodes: CW (central warehouse), AP (airport, capacity: 5 tons) - Demand nodes: DC1, DC2, DC3 (distribution centers) - Critical nodes: H1, H2 (hospitals), FS (fuel station) Resources: $cap_{truck}=4\text{ tons},\tau_{truck}=120\text{ min}$ $cap_{heli}=1\text{ ton},\tau_{heli}=30\text{ min}$ Demand Requirements: - DCs: food/water delivery by $T_{dc}=20:00$ - Hospitals: medicine within $\Delta T_{h}=6$ hours - Fuel Station: refuel by $T_{fs}=12:00$ Critical Deadlines: - Food/Water: All DCs by 8:00 PM - Medicine: Hospitals within 6 hours - Fuel: FS by 12:00 PM - Airport: Clear excess beyond 5 tons immediately Dynamic Disruptions: - Unpredictable donation arrivals - Road blockages requiring rerouting - Emergency hospital demands - Fuel shortage delays Key Planning Requirements: - Resource distribution scheduling - Transportation mode optimization - Delivery prioritization - Airport overflow management - Real-time disruption handling
3.8. P8: Wedding Reunion Planner and Scheduler with Disruptions (WR-dynamic)
@@
3.9. P9: Thanksgiving Dinner Planner and Scheduler with Disruptions (TD-dynamic)
@@
For example, consider James’s flight, which was originally scheduled to arrive in Boston at 1 PM but has been delayed to 4 PM. He learns of this new arrival time at 10 AM EST (the flight’s intended departure from SFO), which offers a three-hour delay for adjustments. However, existing LLM-based solutions do not adapt the plan until 1 PM, thus squandering the opportunity to re-optimize the schedule.
3.10. P10: Global Supply Chain Planner and Scheduler (GSC-static and GSC-dynamic, Table 8)
Table 8. Data Center Construction Problem Statement
GPU Procurement & Shipments Total GPU Target: 100,000 units Vendors: NVIDIA: \$15k/unit, 200k/quarter, 20% maintenance risk/year AMD: \$10k/unit, 150k/quarter, 50% maintenance risk/year Order Timing: All orders on Day 1 Shipment Schedule: Quarterly (e.g., 600k units = Q1, Q2, Q3) Payment Terms: 50% at order, 50% at delivery Natural Disaster Impacts Probabilities: 10% per quarter for Earthquake/Typhoon Delays: 1-month cascading delay Cost Impact: 30% price hike for affected shipment Expedition: Expedite one shipment by 1 month (all shift) Cluster Construction & Infrastructure Definition: 50,000 GPUs/cluster; clusters built concurrently Networking Infrastructure (per cluster) Duration: 1 month Cost: \$25M Dependencies: Starts after GPU, power, cooling readiness Power & Cooling (per cluster) Need: 150 MW + 1M gallons/day cooling Lead Time: Regular: 2 months at \$30M; Expedited: 1 month at \$75M Testing & Certification (per cluster) Normal: 2 months, \$15M Expedited: 1 month, \$45M Options: Expedite any module (150% cost, 50% time) Delay Cost Model Penalty: \$50M revenue + \$10M operations per month Objective Minimize total cost while meeting 15-month deadline Includes procurement, infra, delay penalties, expedition costs
Problem Specification: Table 8 in Appendix A presents a comprehensive problem in data center GPU deployment that captures the complexity of large-scale infrastructure projects. The objective is to complete a 1 million GPU data center in 15 months while minimizing total costs. The problem encompasses procurement decisions between NVIDIA (1515k/unit) and AMD (1010k/unit) GPUs, where each vendor has different maintenance risks (20% vs 50% of unit price over one year) and quarterly shipment capacities.